Radiation makes it impossible to stand close to a hot lava flow. Calculate the rate of heat transfer by radiation from 1.21 m2 of 1276 ºC fresh lava into 37 ºC surroundings, assuming lava’s emissivity is 1.00


1
Expert's answer
2021-03-24T19:44:05-0400

Convert temperature in degree Celsius to Kelvin:

T2=1276+273.15=1549.15 K,

T1=37+273.15=310.15 K.

The net rate of heat radiation is


"\frac{Q}{t}=A\epsilon\sigma(T_2^4-T_1^4),\\\space\\
\frac{Q}{t}=1.21·1·5.67·10^{-8}(1549.15^4-310.15^4)=\\=394500\text{ W}."